# On the connection between Bochner’s theorem on positive definite maps and Choi theorem on complete positivity

### v1

@@ Abstract
In this work, we establish a connection between Bochner’s theorem on positive definite maps and Choi theorem on complete positivity. We begin by defining a convolution product between maps from the contracted semigroup algebra $\mathbb{C}_{0}[S]$ of a semigroup $S$ to an arbitrary associative algebra $\mathcal{A}$. The convolution product makes the space $L(\mathbb{C}_{0}[S],\mathcal{A})$ of linear maps from $\mathbb{C}_{0}[S]$ to $\mathcal{A}$ an associative algebra. We prove that the convolution algebra $L(\mathbb{C}_{0}[S],\mathcal{A})$ and the tensor product algebra $\mathbb{C}_{0}[S]\otimes\mathcal{A}$ are isomorphic. As a consequence, in the specific case of the inverse semigroup of matrix units, we identify the product in the space of maps on the matrix algebras which is preserved by the Choi-Jamiołkowski isomorphism as convolution. Then, by defining the Fourier transform of a map from $\mathbb{C}_{0}[S]$ to $M_{n}(\mathbb{C})$, we derive the Fourier inversion formula when $S$ is a finite inverse semigroup. As a corollary of this formula, we show that in the case of the inverse semigroup of matrix units, the Fourier transformation of a map with respect to the identity representation becomes the Choi matrix of the map and the Fourier inversion formula becomes the Choi inversion formula. Then, by defining the notion of matrix valued positive definite maps, we prove Bochner’s theorem in the context of finite inverse semigroup. It is demonstrated that Bochner’s theorem reduces to Choi theorem on completely positive maps when the inverse semigroup of matrix units is considered. Additionally, the necessary and sufficient condition on a representation $\rho:M_{m}\to M_{d_{\rho}}(\mathbb{C})$ such that the Complete positivity vs. positivity correspondence holds between a linear map $\Phi:M_{m}(\mathbb{C})\to M_{n}(\mathbb{C})$ and its Fourier Transform $\widehat{\Phi}(\rho)$ is obtained.
I Introduction
@@
In the context of quantum theory, linear maps between spaces of operators are essential for describing the evolution of quantum states. Specifically, a quantum state is represented by a density operator—a positive semidefinite matrix with unit trace. A physical process is then modeled by a linear map that sends density operators to other density operators. However, not every such map corresponds to a valid physical transformation. A key requirement is that the map must be completely positive and trace-preserving. Such maps are called quantum channels. Complete positivity ensures that even when the system is entangled with another (possibly unknown) system, the overall transformation remains valid. This is crucial because in quantum mechanics, systems are often not isolated. A foundational result that provides a practical criterion for complete positivity is the Choi theorem [1]. It states that a linear map is completely positive if and only if its associated Choi matrix is positive semidefinite. This correspondence makes it easier to verify whether a given map can represent a physical process.
The Fourier transform is a fundamental tool in both physics and mathematics, with widespread applications across numerous domains. The Fourier transform of an integrable function $f(x)$ on the real line is defined as: $\widehat{f}(t):=\int_{-\infty}^{\infty}e^{-ixt}f(x)dx$ for all $t\in\mathbb{R}$ with the inverse transform given by $f(x):=\frac{1}{2\pi}\int_{-\infty}^{\infty}e^{ixt}\widehat{f}(t)dt$ for all $x\in\mathbb{R}$. These definitions can be generalized for complex valued function on a finite group as follows: Let $G$ be a finite group and $f:G\to\mathbb{C}$ be a complex valued function. Then the Fourier transform of $f$ with respect to a representation $\rho:G\to M_{d_{\rho}}(\mathbb{C})$ is defined as $\widehat{f}(\rho):=\sum_{g\in G}\rho(g)f(g)$ with the inverse transform given by $f(g)=\frac{1}{|G|}\sum_{\rho\in\mathrm{Irr}(G)}d_{\rho}\operatorname{tr}(\rho(g^{-1})\widehat{f}(\rho))$, where $\mathrm{Irr}(G)$ denotes a complete set of inequivalent irreducible representations of $G$. A function $f:G\to\mathbb{C}$ is called positive definite if $\sum_{g,t\in G}f(g^{-1}t)c_{t}\overline{c}_{g}\geq 0$ for any set of complex numbers $\{c_{g}\}_{g\in G}$. Positive definite functions are completely characterized by Bochner’s theorem which states that a function $f:G\to\mathbb{C}$ is positive definite if and only if its Fourier transforms $\widehat{f}(\rho))$ are positive semidefinite for $\rho\in\mathrm{IrrU}(G)$, where $\mathrm{IrrU}(G)$ represents a complete set of inequivalent irreducible unitary representations of $G$. For a map $\Phi:G\to M_{n}(\mathbb{C})$, its Fourier transform with respect to a representation $\rho$ can be defined as $\widehat{\Phi}(\rho):=\sum_{g\in G}\rho(g)\otimes\Phi(g)$. A map $\Phi:G\to M_{n}(\mathbb{C})$ is called positive definite if the matrix $\Phi(g^{-1}g^{\prime})]$ is positive semidefinite. In this case Bochner’s theorem states that a map $\Phi:G\to M_{n}(\mathbb{C})$ is positive definite if and only if its Fourier transforms $\widehat{\Phi}(\rho))$ are positive semidefinite for $\rho\in\mathrm{IrrU}(G)$ (details are provided in the main text).
Now it can be observed that the Choi matrix of a linear map $\Phi:M_{m}(\mathbb{C})\to M_{n}(\mathbb{C})$, defined as $C_{\Phi}:=\sum_{i,j}e_{ij}\otimes\Phi(e_{ij})$, where $\{e_{ij}\}_{i,j=1}^{m}$ denotes the standard matrix units of $M_{m}(\mathbb{C})$, can be interpreted as the Fourier transform of $\Phi$. In particular, the expression $C_{\Phi}:=\sum_{i,j}\mathrm{id}_{m}(e_{ij})\otimes\Phi(e_{ij})$ closely resembles the standard definition of the Fourier transform of maps on finite groups, as discussed above. Moreover, the convolution-like product between two linear maps $\Phi,\Phi^{\prime}:M_{m}(\mathbb{C})\to M_{n}(\mathbb{C})$, defined by $(\Phi*\Phi^{\prime})(e_{ij}):=\sum_{k}\Phi(e_{ik})\Phi^{\prime}(e_{kj})$, is preserved under the Choi–Jamiołkowski isomorphism; that is, $C_{\Phi*\Phi^{\prime}}=C_{\Phi}C_{\Phi^{\prime}}$ [2]. This equation resembles the well-known convolution theorem, which states that the Fourier transform of a convolution is the product of the individual Fourier transforms. These structural parallels motivate us to define a Fourier transform for matrix-valued maps on the contracted algebra, prove Bochner’s theorem in the context of finite inverse semigroups, and establish a connection with Choi theorem on completely positive maps.
I.1 Our Contributions and Overview of our Main Results
@@
With the aim to establish a connection between Bochner’s theorem and Choi theorem, we begin by defining a convolution product between maps from the contracted semigroup algebra $\mathbb{C}_{0}[S]$ of a semigroup $S$ to an arbitrary associative algebra $\mathcal{A}$. The convolution product makes the space $L(\mathbb{C}_{0}[S],\mathcal{A})$ of linear maps from $\mathbb{C}_{0}[S]$ to $\mathcal{A}$ an associative algebra. Then we proved that the convolution algebra $L(\mathbb{C}_{0}[S],\mathcal{A})$ and the tensor product algebra $\mathbb{C}_{0}[S]\otimes\mathcal{A}$ are isomorphic (see Theorem (1). As a consequence of Theorem (1), in the specific case of matrix unit semigroup, we have identified that the product in the space of maps on the matrix algebras which is preserved by the Choi-Jamiołkowski isomorphism is convolution. Then defining the Fourier transform of a map from $\mathbb{C}_{0}[S]$ to $M_{n}(\mathbb{C})$, we have derived the Fourier inversion formula when $S$ is a finite inverse semigroup in Theorem (2) which is one of our main results. As a corollary of Theorem (2) we have shown that in the case of the inverse semigroup of matrix units, the Fourier transformation of a map with respect to the identity representation becomes the Choi matrix of the map and the Fourier inversion formula becomes the Choi inversion formula [see Corollary (1)]. Additionally, We also have derived the Plancherel formula in the context of inverse semigroup [see Theorem (3)].
@@
Then, by defining the notion of matrix valued positive definite maps, we prove Bochner’s theorem in the context of finite inverse semigroup in Theorem (4) which is one of our main results. While proving Bochner’s theorem for finite inverse semigroup, we have provided two useful alternative characterization of positive definite maps on contracted semigroup algebras in Proposition (3) and Proposition (4). Also, we have proved Schur’s orthogonality relation in the context of finite inverse semigroup in Proposition (5) which played a crucial role in the proof of Bochner’s theorem. Then in subsection (III.4) we have shown how Bochner’s theorem reduces to Choi theorem on completely positive maps. Finally, in Theorem (6), we have derived the necessary and sufficient condition on a representation $\rho:M_{m}\to M_{d_{\rho}}(\mathbb{C})$ such that the Complete positivity vs. positivity correspondence holds between a linear map $\Phi:M_{m}(\mathbb{C})\to M_{n}(\mathbb{C})$ and its Fourier Transform $\widehat{\Phi}(\rho)$.
I.2 Structure of the paper
@@
The structure of the paper is as follows: In Section II, we provide basic ideas of a semigroup, inverse semigroup and their contracted algebras. Section III contains our main results. It is divided into several subsections. In Subsection III.1, we define and discuss various properties of Fourier transform of maps on the contracted semigroup algebra of a finite semigroup. In Subsection III.2, we derive the Fourier inversion formula for finite inverse semigroup and show that in the case of matrix unit semigroup, the formula reduces to the Choi inversion formula. Additionally, Plancherel formula in the context of finite inverse semigroup is also derived. By defining the notion of matrix valued positive definite maps, we prove Bochner’s theorem in the context of finite inverse semigroup in Subsection III.3. Also we provide two useful alternative characterizations of positive definite maps on the contracted semigroup algebra of a finite inverse semigroup. Here, we prove Schur’s orthogonality relation in the context of finite inverse semigroup which played a crucial role in the proof of Bochner’s theorem. Then, in subsection III.4, we show how Bochner’s theorem reduces to Choi theorem on completely positive maps. In Subsection III.5, we derive the necessary and sufficient condition on a representation $\rho:M_{m}\to M_{d_{\rho}}(\mathbb{C})$ such that the Complete positivity vs. positivity correspondence holds between a linear map $\Phi:M_{m}(\mathbb{C})\to M_{n}(\mathbb{C})$ and its Fourier Transform $\widehat{\Phi}(\rho)$. Finally, we conclude our work in Section IV.
II Preliminaries
@@
In this section, we briefly discuss some basics of finite inverse semigroup, its useful properties and representations. Also we discuss Choi theorem on completely positive maps briefly.
II.1 Semigroup, Inverse Semigroup and their Algebras
@@
A semigroup is a pair $(S,\cdot)$ of a non-empty set $S$ and a binary operation $\cdot:S\times S\to S$ which is associative, i.e., $x\cdot(y\cdot z)=(x\cdot y)\cdot z$ for all $x,y,z\in S$. Throughout the paper, we adopt the notation that $x\cdot y=xy$.
Definition.
@@
(Semigroup Algebra) Let $S$ be a semigroup. A semigroup-algebra of $S$ over the complex field $\mathbb{C}$ is an algebra $\mathbb{C}[S]$ with a basis which is isomorphic to the semigroup $S$ under the multiplication rule in $\mathbb{C}[S]$. With the abuse of notation, we will use $S$ itself to denote the basis for $\mathbb{C}[S]$.
Definition.
@@
A contracted algebra $\mathbb{C}_{0}[S]$ of a semigroup $S$ is an algebra over $\mathbb{C}$, which have a basis $\mathscr{B}$ such that $\mathscr{B}\cup\{0\}$ is isomorphic to $S$, where 0 represents the zero of the contracted algebra.
@@
With the abuse of notation, we will use $S\setminus\{0\}$ (or sometimes $S\setminus z$) to denote a basis for $\mathbb{C}_{0}[S]$. In the following, we provide a brief description of the basic concepts and properties of an inverse semigroup. The concept of an inverse semigroup provides the notion of inverse without having the notion of an identity element unlike a group.
Definition A.
@@
(Inverse Semigroup) A semigroup $S$ is called an inverse semigroup if for every element $s\in S$ there exists a unique element denoted by $s^{-1}$ such that $ss^{-1}s=s$ and $s^{-1}ss^{-1}=s^{-1}$.
@@
It trivially follows from the definition that $(s^{-1})^{-1}=s$. An element $p\in S$ is called an idempotent element if $p^{2}=p$. Now we list some elementary properties that hold in an inverse semigroup.
Proposition.
Let $S$ be an inverse semigroup. Then the following properties holds.
@@
    For any element $s\in S$, the elements $ss^{-1}$ and $s^{-1}s$ are idempotents.
@@
    Let $e$ and $f$ be two idempotents in $S$. Then $ef$ is also an idempotent and $ef=fe$. i.e., the idempotents commutes.
@@
II.1.1 Partial ordering and Möbius Inversion
Let $S$ be a finite partially ordered set and $R$ be a commutative unital ring. The incidence algebra denoted by $R[S\times S]$ is an algebra of all the functions $f:S\times S\to R$ such that whenever $x\in S$ and $y\in S$ are not compatible i.e., they are not related by the partial ordering in $S$ ($x\nleq y$ or $y\nleq x$), we have $f(x,y)=0$. The multiplication between $f,h\in R[S\times S]$ is the convolution defined as follows:

### v2

# arXiv:2509.02529v2[math-ph] 02 Mar 2026

@@ Abstract
Bochner’s theorem characterizes positive definite functions on groups through the positivity of their Fourier transforms and plays a fundamental role in Harmonic analysis. While Bochner-type results are known for certain classes of semigroups, they typically differ from the group theoretic formulations and do not retain the same level of simplicity and generality. In this work, we prove a Bochner-type theorem for finite inverse semigroups at the level of matrix valued linear maps on the contracted algebras of the semigroups. Using the intrinsic partial order of inverse semigroups, positivity naturally arises through a Möbius-transformed map. Our main result characterizes the positive definiteness of the Möbius transformed map in terms of the positivity of the Fourier transform of the original map with respect to a complete family of inequivalent irreducible representations of the contracted algebra induced by irreducible unitary representations of the maximal subgroups of the inverse semigroup. The proof relies on Fourier inversion formula, Schur orthogonality relations, and alternative characterizations of positive definite maps, all established here in the setting of finite inverse semigroups. As a special case, we show that for the inverse semigroup of matrix units, Bochner’s theorem reduces exactly to Choi’s characterization of completely positive maps.
keywords:
Positive definite maps, Bochner’s theorem, Inverse semigroups, Fourier transforms, Choi matrix, Completely positive maps
1 Introduction
Positive definite objects and Fourier analysis play a fundamental role in mathematics and physics. In particular, Bochner’s theorem provides a complete characterization of positive definite functions on groups in terms of the positivity of their Fourier transforms [1, 2, 3]. This result underlines large parts of harmonic analysis, representation theory, and probability theory, and has far reaching consequences in both classical and quantum setting [4, 5].
Bochner-type theorems have also been studied in various semigroup settings, particularly for abelian semigroups with involution and topological semigroups [6]. Although these formulations capture important aspects of positivity, they differ substantially from the group setting, both in their analytic framework and in the role played by representation theory, and do not hold with the same simplicity and generality even for locally compact Hausdorff involution semigroups [7]. Bochner-type results have also been obtained in specific scenarios of finite-dimensional conelike semigroups [8], discrete commutative semigroups with identity and an involution [9, 10], foundation topological semigroups [11].
Inverse semigroups [12] provide a natural algebraic framework for describing partial symmetries and arise prominently in the study of groupoids and operator algebraic constructions. Their representation theory is well understood and admits a nice structure via maximal subgroups and induced representations[13, 14]. However, a Bochner-type characterization of positive definite operator-valued maps on finite inverse semigroups does not appear to have been previously developed to the best of our knowledge.
One of the main objectives of this work is to establish a Bochner-type theorem for finite inverse semigroups. More precisely, we formulate and prove a Bochner-type theorem for matrix valued maps on the contracted algebra of a finite inverse semigroup. Due to the intrinsic partial order of inverse semigroups, positivity arises at the level of Möbius transformed map. Our main result characterizes the positive definiteness of the Möbius transformed map in terms of the positivity of the Fourier transform of the original map with respect to a complete family of inequivalent irreducible representations of the contracted algebra induced by a complete family of inequivalent irreducible unitary representations of the maximal subgroups of the inverse semigroup. The proof relies on Fourier inversion formula, Schur orthogonality relations, and alternative characterizations of positive definite maps, all developed here in the setting of finite inverse semigroups. This framework provides a genuine extension of Bochner’s theorem beyond groups, while retaining structural similarities.
@@ 1 Introduction
In quantum theory, linear maps between operator algebras encode the dynamics of physical systems. A quantum state is represented by a density operator—a positive semidefinite matrix with unit trace. A physical process is then modeled by a linear map that sends density operators to other density operators. However, not every such map corresponds to a valid physical transformation. A key requirement is that the map must be completely positive and trace-preserving. Such maps are called quantum channels. Complete positivity ensures that even when the system is entangled with another (possibly unknown) system, the overall transformation remains valid. This is crucial because in quantum mechanics, systems are often not isolated. A foundational result that provides a practical criterion for complete positivity is the Choi theorem [15]. It states that a linear map is completely positive if and only if its associated Choi matrix is positive semidefinite. Choi’s theorem and the Choi matrix are essential tools for the mathematical analysis of quantum channels and superchannels [16, 17, 18, 19, 20, 21, 22, 23, 24].
At first sight, Bochner’s theorem and Choi’s theorem arise in rather different settings: the former concerns positive definite functions or maps on groups and their Fourier transforms, while the latter provides matrix positivity criterion for completely positivity of linear maps between operator algebras. Nonetheless, both results are, from a broader perspective, statements relating positivity in an “original domain” to positivity in a “transformed domain”. This observation leads to the following question: Can Choi’s theorem on complete positivity be understood as a special case of Bochner-type theorem? Our formulation of the Bochner-type theorem answers this question in affirmative. In particular, this serves as an important and illuminating special case of our general results when the inverse semigroup of matrix units is considered. The full matrix algebra over the complex field is a contracted semigroup algebra of the inverse semigroup of matrix units [25]. In this setting, the Fourier transform of a linear map on the matrix algebra coincides exactly with the Choi matrix of the map, and the Bochner-type theorem reduces to Choi’s theorem on complete positivity. From this perspective, Choi’s theorem emerges as a special instance of a more general Bochner-type result, rather than as an isolated criterion.
1.1 Our Contributions and Overview of the Main Results
@@
The main goal of this work is to establish a Bochner’s theorem for finite inverse semigroups and to recover Choi’s theorem on complete positivity as a special case. We begin by defining a convolution product between maps from the contracted semigroup algebra $\mathbb{C}_{0}[S]$ of a semigroup $S$ to an arbitrary associative algebra $\mathcal{A}$. The convolution product makes the space $L(\mathbb{C}_{0}[S],\mathcal{A})$ of linear maps from $\mathbb{C}_{0}[S]$ to $\mathcal{A}$ an associative algebra. Then we show that the convolution algebra $L(\mathbb{C}_{0}[S],\mathcal{A})$ and the tensor product algebra $\mathbb{C}_{0}[S]\otimes\mathcal{A}$ are isomorphic [see Theorem 1]. As a consequence of Theorem 1, in the specific case of the matrix unit semigroup, we identify the product in the space of maps on the matrix algebras that is preserved by the Choi-Jamiołkowski isomorphism as convolution. Then, defining the Fourier transform of a map from $\mathbb{C}_{0}[S]$ to $M_{n}(\mathbb{C})$, we derive a Fourier inversion formula when $S$ is a finite inverse semigroup [ see Theorem 2], which is one of our main results. As a corollary of Theorem 2, we show that for the inverse semigroup of matrix units, the Fourier transform of a map with respect to the identity representation coincides with the Choi matrix of the map, and the Fourier inversion formula reduces to the Choi inversion formula [see Corollary 1]. Additionally, We derive a Plancherel formula in the finite inverse semigroup setting [see Theorem 3].
@@
Then, by introducing the notion of matrix valued positive definite maps on contracted semigroup algebras, we prove a Bochner’s theorem for finite inverse semigroups in Theorem 4, which is the central result of this work. The proof relies on Fourier inversion formula [see Theorem 2], Schur orthogonality relations [see Proposition 5], and alternative characterizations of positive definite maps [see Proposition 3, and Proposition 4], all developed here in the setting of finite inverse semigroups. Then, in Section 3.4, we show that Bochner’s theorem reduces to Choi theorem on completely positive maps when inverse semigroup of matrix units are considered. Finally, in Theorem 6, we derive a necessary and sufficient condition on a representation $\rho:M_{m}(\mathbb{C})\to M_{d_{\rho}}(\mathbb{C})$ such that the Complete positivity vs. positivity correspondence holds between a linear map $\Phi:M_{m}(\mathbb{C})\to M_{n}(\mathbb{C})$ and its Fourier Transform $\widehat{\Phi}(\rho)$.
1.2 Structure of the paper
@@
The structure of the paper is as follows. In Section 2, we review basic notions of semigroups, inverse semigroups, their contracted algebras, and the corresponding representation theory. Section 3 contains the main results of the paper and is divided into several subsections. In Section 3.1, we define the Fourier transform of maps on the contracted algebra of a finite inverse semigroup and discuss its various properties. In Section 3.2, we derive a Fourier inversion formula for finite inverse semigroups and show that in the case of inverse semigroup of matrix units, this formula reduces to the Choi inversion formula. A Plancherel formula is also derived in the finite inverse semigroup setting. In Section 3.3, we prove a Bochner’s theorem for finite inverse semigroups. Two alternative characterizations of positive definite maps on contracted algebras of finite inverse semigroups are provided, and a Schur orthogonality relation in the setting of finite inverse semigroups is proved, which plays a crucial role in the proof of Bochner’s theorem. Section 3.4 is devoted to the finite dimensional matrix algebra case, where we show that Bochner’s theorem reduces to Choi’s characterization of completely positive maps. In Section 3.5, we derive a necessary and sufficient condition on a representation $\rho:M_{m}(\mathbb{C})\to M_{d_{\rho}}(\mathbb{C})$ such that the Complete positivity vs. positivity correspondence holds between a linear map $\Phi:M_{m}(\mathbb{C})\to M_{n}(\mathbb{C})$ and its Fourier Transform $\widehat{\Phi}(\rho)$. Finally, Section 4 contains concluding remarks and perspectives.
2 Preliminaries
@@
In this section, we provide a brief overview of the fundamental concepts of finite inverse semigroups along with their key properties and representations. We also present a concise discussion of Choi’s theorem on completely positive maps.
2.1 Semigroup, Inverse Semigroup and their Algebras: Structure and Representations
@@
A semigroup is a pair $(S,\cdot)$ of a non-empty set $S$ and a binary operation $\cdot:S\times S\to S$ which is associative, i.e., $x\cdot(y\cdot z)=(x\cdot y)\cdot z$ for all $x,y,z\in S$. Throughout the paper, we adopt the notation $x\cdot y=xy$.
Definition (Semigroup Algebra [26]).
@@
Let $S$ be a semigroup. A semigroup algebra of $S$ over the complex field $\mathbb{C}$ is an algebra $\mathbb{C}[S]$ with a basis which is isomorphic to the semigroup $S$.
By the abuse of notation, we use $S$ itself to denote the basis for $\mathbb{C}[S]$.
Definition (Contracted Algebra [26]).
@@
A contracted algebra $\mathbb{C}_{0}[S]$ of a semigroup $S$ is an algebra over the complex field $\mathbb{C}$ with a basis $\mathscr{B}$ such that $\mathscr{B}\cup\{0\}$ is isomorphic to $S$, where 0 denotes the zero of the contracted algebra.
@@
By the abuse of notation, we use $S\setminus\{0\}$ (or sometimes $S\setminus\{z\}$) to denote a basis of $\mathbb{C}_{0}[S]$. In the following, we provide a brief description of the basic concepts and properties of an inverse semigroup. Unlike a group, an inverse semigroup generalizes the notion of inverse without requiring the presence of an identity element.
Definition A (Inverse Semigroup [26, 12]).
@@
A semigroup $S$ is called an inverse semigroup if, for every element $s\in S$, there exists a unique element $s^{-1}$ such that $ss^{-1}s=s$ and $s^{-1}ss^{-1}=s^{-1}$.
@@
It immediately follows from the definition that $(s^{-1})^{-1}=s$. An element $p\in S$ is called idempotent if $p^{2}=p$. We now list some elementary properties that hold in an inverse semigroup.
Proposition ([26, 12]).
Let $S$ be an inverse semigroup. Then the following properties hold.
@@
    For any element $s\in S$, the elements $ss^{-1}$ and $s^{-1}s$ are idempotent.
@@
    Let $e$ and $f$ be two idempotents in $S$. Then $ef$ is also idempotent, and $ef=fe$, i.e., the idempotents commute.
@@
2.1.1 Partial Ordering and Möbius Inversion
Let $S$ be a finite partially ordered set and $R$ a commutative unital ring. The incidence algebra of $S$ over $R$, denoted by $R[S\times S]$, is an algebra of all the functions $f:S\times S\to R$ such that
$f(x,y)=0$ (1)
whenever $x,y\in S$ are incomparable i.e., $x\nleq y$. The multiplication between $f,h\in R[S\times S]$ is given by the convolution defined as follows [27]: